Find the mean
Find the median
Find the mode
Find the upper and lower quartiles
Find the range
Find the variance
Find the standard deviation
79.6,94.8,94.2,86,83.6,82,89.4,76.4,85,92.9,95.6,90.6,91.7,87.9,85.7,91.7,68,88,78.9,98.2,86.3,96
I am really having a hard time figuring this out, any help would be greatly appreciated!
THanks!

Answer by oberobic(2304)

The first thing to do is to get the data sorted into ascending order. You will need them sorted to figure out the median.
68.0
76.4
78.9
79.6
82.0
83.6
85.0
85.7
86.0
86.3
87.9
88.0
89.4
90.6
91.7
91.7
92.9
94.2
94.8
95.6
96.0
98.2
Add them up = 1922.5
There are 22 items in the list so the mean (arithmetic mean) is the total divided by the number of items.
1922.5/22 = 87.38636364
The median is the "middle" of a sorted list. In this case the middle is between the 11th and the 12th items in the list.
#11 = 87.9
#12 = 88.0
So, the median splits the difference
(87.9+88.0)/2 = 87.95
The mode is the most frequently occurring data value.
The only number that occurs more than once is 91.7, which occurs twice.
So the mode = 91.7
The lower quartile is where 25% of value are less than or equal. The upper quartile is where 75% are less than or equal.
The minimum value is 68.0
The maximum value is 98.2
So the range = max - min = 98.2 - 68 = 30.2
The standard deviation depends on finding the mean first.
Then you subtract the mean from each of the values.
Then you square each of these numbers and sum them.
Then take the square root.
(If you calculated the mean correctly, then the sum of each of the original values minus the mean is zero...which is not very useful. But by squaring these differences, then you have something that you can use to describe dispersion.)
.
ItemValueValue-MeanSquared
1 68.0 -19.4 375.83110
2 76.4 -11.0 120.70019
3 78.9 -8.5 72.01837
4 79.6 -7.8 60.62746
5 82.0 -5.4 29.01291
6 83.6 -3.8 14.33655
7 85.0 -2.4 5.69473
8 85.7 -1.7 2.84382
9 86.0 -1.4 1.92200
10 86.3 -1.1 1.18019
11 87.9 0.5 0.26382
12 88.0 0.6 0.37655
13 89.4 2.0 4.05473
14 90.6 3.2 10.32746
15 91.7 4.3 18.60746
16 91.7 4.3 18.60746
17 92.9 5.5 30.40019
18 94.2 6.8 46.42564
19 94.8 7.4 54.96200
20 95.6 8.2 67.46382
21 96.0 8.6 74.19473
22 98.2 10.8 116.93473
Sum 1922.5 0.0 1126.8

Mean = 1922.5/22 = 87.3863
Variance = Sum of Squared Values divided by # items minus 1
Variance = 1126.8/21 = 53.65647186
Std Dev = Sqrt(Variance) = Sqrt(1126.8) = 33.56763
The quartiles are 83.95 and 92.6
Done